# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v5

@@ 3.2 Negative sample re-weighting strategy
To address the second challenge of the class imbalance issue, we employ a negative sample re-weighting strategy to minority AUs. This also contributes to hard negative sample mining. Inspired by HCL [30], we apply the following weighting scheme to each negative sample $x_{j}^{-}$:
@@
Above equation follows the Boltzmann distribution. As the value of $\beta$ increases, the entropy of the distribution decreases significantly [31], leading to a sharper distribution in the region of high similarity. This has two consequences: 1. It imposes larger penalties on hard negative samples that are in close proximity to the anchor data point, enabling the model to learn more precise decision boundaries from subtle feature differences when AUs activate. 2. It controls the relative contribution of negative samples to the parameter updates, providing a means to address class imbalance issues.
The primary distinction between our weighting method and HCL [30] lies in the fact that Eq. (3) does not divide by the mean as a normalization factor, enabling a more effective solution to the class imbalance issue in the field of AU detection. Specifically, when anchor data point $x$ belongs to the majority class, its positive sample belongs to the majority class, while its negative sample belongs to the minority class. In this case, we aim to have smaller gradient magnitudes for the positive samples (majority class) to update fewer parameters, while the negative samples (minority class) should have larger gradient magnitudes to update more parameters. To achieve this, we begin to set $\beta\geq 1$, which increases the impact of parameter updates caused by negative samples (minority class). On the other hand, when $x$ belongs to minority class, we begin to set $\beta\leq 1$ to reduce the influence of parameter updates caused by majority class samples. By adjusting the value of $\beta$, we can control the relative importance of positive and negative samples during the parameter updating, ensuring that the model pays more attention to minority class samples when necessary and balances the gradient magnitudes accordingly. This helps to balance the representation of different AUs in training set and improves the model ability to precisely classify all AU samples. If the mean were used as a normalization factor, the precise control over the parameter updates for majority and minority classes through the setting of $\beta$ would become unfeasible. This is because changes in $\beta$ would concurrently alter the mean, rendering the subsequent increase or decrease in weights after division by the mean uncertain. Therefore, the omission of division by the mean in Eq. (3), motivated by considerations of class imbalance, distinguishes our work from HCL.
Figure 3: Illustration of our proposed positive sample sampling strategy.
@@ 3.3.1 Supervised Signal
One common technique to address the issue of false data is label smoothing, where a small amount of uncertainty is incorporated into the training process by replacing one-hot labels with smoothed probability distributions. However, it is challenging to directly apply label smoothing to contrastive learning as the computation of contrastive loss does not involve explicit labels. To overcome this challenge, we turn to the memory effect of neural networks. Previous research has shown that neural networks tend to first fit clean samples and then memorize the noise in incorrectly labeled samples [32]. This memory effect is independent of the choice of neural network architecture. The positive samples that are easier fit are those that are closer to the anchor point and have higher similarity. Inspired by this observation, for a given sample $x$ and M positive samples ${x}_{i=1}^{M}$ with noisy labels belonging to the same class, we calculate the similarity scores between the representation of sample $x$ and the representations of all M positive samples. We then select the positive feature representation with the highest similarity to sample $x$:
@@ 4.1 Experiment Setting
4.1.1 Datasets
In our experiments, we extensively utilize three well-established datasets in constrained scenarios for AU detection, namely BP4D [34], DISFA [35], GFT [36], and one well-established dataset in unconstrained scenarios, namely Aff-Wild2 [37]. Our primary focus is on assessing frame-level AU detection, as opposed to datasets that solely provide video-level annotations.
BP4D [34]: This dataset comprises 41 participants, including 23 females and 18 males. Each individual is featured in 8 sessions, recorded using both 2D and 3D videos. The dataset encompasses approximately 140,000 frames, each annotated with AU labels indicating occurrence or non-occurrence. Additionally, each frame is meticulously annotated with 68 facial landmarks. For BP4D dataset, we employ a three-fold cross-validation methodology for rigorous evaluation.
DISFA [35]: This dataset is composed of 27 videos, recorded from 15 men and 12 women. Each video comprises 4,845 frames, resulting in a total of 130,788 images. Frames in this dataset are annotated with 68 facial landmarks and AU intensities ranging from 0 to 5. Frames with intensities equal to or greater than 2 are considered positive, while others are classified as negative. We also employ the same three-fold cross-validation methodology as BP4D dataset for rigorous evaluation of DISFA dataset.
GFT [36]: This dataset comprises 96 participants, organized into 32 groups of three individuals, engaged in unscripted conversations. Each participant is recorded via video, with the majority of frames capturing moderate out-of-plane poses. Annotations for each frame include 10 AUs, alongside 68 facial landmarks. In accordance with the standard training and testing partitions, our training set consists of 78 participants, amounting to approximately 108,000 frames, while the testing set includes 18 participants, totaling around 24,600 frames.
Aff-Wild2 [37]: This dataset is an extensive in-the-wild facial expression dataset, with videos sourced from YouTube, encompassing a broad range of ages, ethnicities, professions, emotions, poses, lighting conditions, and occlusions. This dataset comprises 305 videos, totaling approximately 1,390,000 frames for the training set, and 105 videos, with around 440,000 frames for the validation set. Each frame is annotated with 12 AUs and 68 facial landmark annotations. Following the methodology of Zhang et al. [38], we train our AUNCE on the training set and evaluate it on the validation set.
@@ 4.1.2 Implementation Details
Initially, the 68 facial landmarks undergo a conversion into 49 facial internal landmarks, excluding those associated with the facial contour. Subsequently, the images undergo augmentation through similarity transformation, encompassing in-plane rotation, uniform scaling, and translation, utilizing the 49 facial landmarks. The resulting images are standardized to a resolution of 256$\times$256$\times$3 pixels and then randomly cropped into 224$\times$224$\times$3 pixels, with the addition of a random horizontal flip.
For the training process, the encoder is conducted on one NVIDIA GeForce RTX 3090 GPU with 24GB for one day and implemented by PyTorch using an AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999 and weight decay of $10^{-6}$. We set a learning rate of $10^{-5}$ and 60 epochs. Due to our discriminative contrastive learning framework that sample pairs are selected in a same batch, so the batch size has a strong influence on the experimental results. In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 12 for BP4D and 8 for DISFA respectively, and the temperature coefficient $\tau$ is 0.5. The analysis for $\beta$ and $p_{1}$, $p_{2}$, $p_{3}$ will introduce in Section. 4.4.2.
In assessing the acquired representations, we adhere to the established linear evaluation protocol, as commonly employed in previous studies [12, 10]. This involves training a linear classifier atop the fixed base network, and the highest test accuracy across our models serves as a surrogate measure for the quality of the learned representations.
@@ 4.1.3 Evaluation Metrics
[…]
@@
AlexNet [49] 2012 NIPS 44 46 2 73 72 82 5 19 43 42 42.8
LSVM [50] 2008 JMLR 38 32 13 67 64 78 15 29 49 44 42.9
TCAE [51] 2019 CVPR 43.9 49.5 6.3 71.0 76.2 79.5 10.7 28.5 34.5 41.7 44.2
CPAM [52] 2020 TBIOM 43.7 44.9 19.8 74.6 76.5 79.8 50.0 33.9 16.8 12.9 45.3
@@ 4.2.1 Evaluation on BP4D
Methods in the first group are based on feature extraction of regional area around AUs, including EAC-Net [4], ROI [3] and ARL [5]. Our approach achieves a notable improvement, with an increase in F1-score ranging from 5.0% to 10.2% over these methods. This advancement is primarily due to early feature extraction methods neglecting critical correlation information among AUs and failing to specifically address the class imbalance issue for each AU. Furthermore, the precision of the encoders utilized by these methods is generally lower.
@@
Methods in the fourth group focus on capturing correlational information among AUs, including MCM [43], AAR [45], ISTR [46], HMP-PS [44] and MEGraph [33]. Our approach exhibits a performance advantage of approximately 0.6%-4.7% in F1-score. This underscores the efficacy of our supervised discriminative contrastive learning paradigm in effectively capturing differential information between AUs from each sample pair instead of pixel-level information of the entire face associated with AU. It is noteworthy that the encoder used in our method is only the same as the network MEGraph used for the first stage of training, with reduced parameters by nearly 7M. This highlights learning only the difference information among features is more efficient than learning the whole pixel-level information.
Methods in the last group are based on contrastive learning, including SIMCLR [10], MoCo [17], CLP [13], BTFM [47], SupHCL [28], KSRL [24] and CLEF [48]. When compared to these methods, our approach demonstrates a performance advantage of approximately 0.2% to 16.3% in F1-score. This improvement is attributed to the efficacy of our proposed negative sample re-weighting strategy and positive sample sampling strategy, specifically tailored for the AU detection task using static images. In addition, compared to contrastive learning methods based on weighted cross-entropy loss, our approach significantly reduces the extraction of pixel-level redundant information.
@@ 4.2.2 Evaluation on DISFA
Comparable outcomes are evident in the DISFA dataset, as outlined in Table 3, where AUNCE attains the highest overall performance in AU detection. Furthermore, a comparison between Table 2 and Table 3 reveals that recent methods such as SEV-Net, MEGraph, and BTFM demonstrate strong performance on the BP4D dataset but achieve only mediocre results on the DISFA dataset. This disparity suggests their limited generalizability. The more pronounced class imbalance issue in DISFA, as shown in Table 1, likely contributes to this performance gap. In contrast, AUNCE consistently delivers robust performance on both BP4D and DISFA datasets and achieves more stable results across various AUs due to our proposed negative sample re-weighting strategy.
@@ 4.2.3 Evaluation on GFT
Table 4 presents the F1-score results on the GFT dataset, where our AUNCE method demonstrates significantly superior performance compared to previous approaches. Unlike BP4D and DISFA, which primarily consist of near-frontal facial images, GFT includes many images captured under out-of-plane poses. In this challenging scenario, other methods such as EAC-Net, ARL and CLP so on struggle to perform well. In contrast, AUNCE achieves a commendable average F1-score of 63.0%.
@@ 4.2.4 Evaluation on Aff-Wild2
Table 6: Comparison of the weighted cross-entropy (WCE) loss and AUNCE loss with different encoders on BP4D dataset.
Backbone Source Loss F1-Score(%) Accuracy(%)
ResNet18 [54] 2016CVPR WCE 52.3 73.8
AUNCE $\textbf{54.9}^{{\color[rgb]{1,0,0}2.6\uparrow}}$ $\textbf{74.3}^{{\color[rgb]{1,0,0}0.5\uparrow}}$
ResNet50 [54] 2016CVPR WCE 56.2 75.5
AUNCE $\textbf{59.7}^{{\color[rgb]{1,0,0}3.5\uparrow}}$ $\textbf{77.0}^{{\color[rgb]{1,0,0}1.5\uparrow}}$
ROI [3] 2017CVPR WCE 56.4 75.9
AUNCE $\textbf{58.8}^{{\color[rgb]{1,0,0}2.4\uparrow}}$ $\textbf{76.7}^{{\color[rgb]{1,0,0}0.8\uparrow}}$
ARL [5] 2019TAC WCE 61.1 78.2
AUNCE $\textbf{62.9}^{{\color[rgb]{1,0,0}1.8\uparrow}}$ $\textbf{79.1}^{{\color[rgb]{1,0,0}0.9\uparrow}}$

### v6

@@ 3.3 Negative sample re-weighting strategy
To address the second challenge posed by class imbalance, we implement a negative sample re-weighting strategy specifically targeting the anchor data points. This also contributes to hard negative sample mining. Inspired by HCL [25], we apply the following weighting scheme to each negative sample $x_{j}^{-}$:
@@
Eq. 7 follows the Boltzmann distribution. As the value of $\beta$ increases, the entropy of the distribution decreases significantly, leading to a sharper distribution in the region of high similarity. This has two consequences: 1. It imposes larger penalties on hard negative samples that are in close proximity to the anchor data point, enabling the model to learn more precise decision boundaries from subtle feature differences when AUs activate. 2. It controls the relative contribution of negative samples to the parameter updates, providing a means to address class imbalance issues.
The key difference between our weighting method and HCL [25] lies in the absence of mean normalization in Eq. (3), allowing for more effective handling of class imbalance in AU detection. Specifically, when the anchor sample $x$ belongs to the majority class, its positive samples (majority class) receive smaller gradient magnitudes to limit parameter updates, while its negative samples (minority class) receive larger gradient magnitudes to emphasize their impact. To achieve this, we set $\beta\geq 1$, increasing the influence of updates caused by minority class samples. Conversely, when $x$ belongs to the minority class, we set $\beta\leq 1$, reducing the influence of majority class samples. By tuning $\beta$, we control the relative importance of positive and negative samples, ensuring the model focuses appropriately on minority class samples and balances gradient magnitudes. Including the mean as a normalization factor would undermine this control, as changes in $\beta$ would simultaneously affect the mean, leading to unpredictable weight adjustments. Thus, the omission of mean normalization in Eq. (3) is a deliberate choice to address class imbalance, differentiating our approach from HCL.
@@ 3.4 Positive sample samplying strategy
Figure 3: Illustration of our proposed positive sample sampling strategy. It demonstrates the sampling strategy used to mitigate the impact of noisy labels by integrating self-supervised and supervised signals. Positive samples are categorized into three types and sampled with varying probabilities, enhancing the model’s robustness to variations and errors in the labeled dataset.
@@ 3.4.1 Supervised Signal
One common technique to address the issue of false data is label smoothing, where a small amount of uncertainty is incorporated into the training process by replacing one-hot labels with smoothed probability distributions. However, it is challenging to directly apply label smoothing to contrastive learning as the computation of contrastive loss does not involve explicit labels. To overcome this challenge, we turn to the memory effect of neural networks. Previous research has shown that neural networks tend to first fit clean samples and then memorize the noise in incorrectly labeled samples. This memory effect is independent of the choice of neural network architecture. The positive samples that are easier fit are those that are closer to the anchor point and have higher similarity. Inspired by this observation, for a given sample $x$ and M positive samples ${x}_{i=1}^{M}$ with noisy labels belonging to the same class, we calculate the similarity scores between the representation of sample $x$ and the representations of all M positive samples. We then select the positive feature representation with the highest similarity to sample $x$:
@@ 4.1 Experiment Setting
4.1.1 Datasets
In our experiments, we extensively utilize four well-established datasets in constrained scenarios for AU detection, namely BP4D [27], DISFA [28], BP4D+ [29], GFT [30], and one well-established dataset in unconstrained scenarios, namely Aff-Wild2 [31]. Our primary focus is on assessing frame-level AU detection, as opposed to datasets that solely provide video-level annotations.
BP4D [27]: This dataset includes 41 participants (23 females, 18 males), each appearing in 8 sessions with 2D and 3D video recordings. It contains approximately 140,000 frames, each labeled for AU occurrence and annotated with 68 facial landmarks. A three-fold cross-validation is used for evaluation.
DISFA [28]: Comprising 27 videos from 15 men and 12 women, this dataset includes 130,788 frames annotated with 68 facial landmarks and AU intensities (0-5). Frames with intensities $\geq$ 2 are positive; others are negative. Three-fold cross-validation is applied for evaluation, similar to BP4D.
BP4D+ [29]: Featuring 82 females and 58 males across 10 sessions, this dataset expands upon BP4D in scale and diversity. Of the 10 sessions, 4 provide 197,875 AU-labeled frames annotated with 49 facial landmarks. Following [6], we train on BP4D and test on BP4D+.
GFT [30]: This dataset records 96 participants in 32 unscripted conversational groups, primarily capturing moderate out-of-plane poses. Annotations include 10 AUs and 68 facial landmarks. Training involves 78 participants (108,000 frames), while testing uses 18 participants (24,600 frames) based on standard splits.
Aff-Wild2 [31]: This in-the-wild dataset includes 305 training videos (1.39 million frames) and 105 validation videos (440,000 frames), sourced from YouTube with diverse demographics and conditions. Each frame is annotated with 12 AUs and 68 facial landmarks. We follow [32] by training on the training set and evaluating on the validation set.
@@ 4.1.2 Implementation Details
Following the preprocessing method used in MEGraph [26], the 68 facial landmarks undergo a conversion into 49 facial internal landmarks, excluding those associated with the facial contour. Subsequently, the images undergo augmentation through similarity transformation, encompassing in-plane rotation, uniform scaling, and translation, utilizing the 49 facial landmarks. The resulting images are standardized to a resolution of 256$\times$256$\times$3 pixels and then randomly cropped into 224$\times$224$\times$3 pixels, with the addition of a random horizontal flip.
For the pre-training process, the encoder is conducted on one NVIDIA GeForce RTX 3090 GPU with 24GB for one day and implemented by PyTorch using an AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999 and weight decay of $10^{-6}$. We set a learning rate of 10-5 and 60 epochs. In our discriminative contrastive learning framework, where sample pairs are selected within the same batch, the batch size significantly impacts the experimental results. For consistency, we have set the batch size to 32 in our experiments. In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 12 for BP4D, BP4D+, Aff-Wild2, 10 for GFT, and 8 for DISFA respectively, and the temperature coefficient $\tau$ is 0.5. The analysis for $\beta$ and $p_{1}$, $p_{2}$, $p_{3}$ will introduce in Section. 4.4.1.
For the linear evaluation protocol [9, 13], we extract $n_{au}$ 2048-dimensional feature vectors from the model pre-trained using the AUNCE loss. A linear classifier is then trained on these features using AdamW optimizer with the same settings as the pre-training phase, a batch size of 256 on a single GPU, a learning rate of 10-3, and 100 epochs. The results are obtained through training with the weighted cross-entropy loss (WCE), and the highest test accuracy across our models serves as a surrogate measure for the quality of the learned representations.
@@ 4.1.3 Evaluation Metrics
Aligned with current state-of-the-art methodologies, we evaluate our AU detection method using three frame-based metrics: F1-macro score, F1-micro score, and Accuracy. The F1-macro score is obtained by calculating the F1 score for each AU class individually and then averaging these values, which is a standard metric employed by many existing methods. Inspired by the work of Hinduja et al. [33], we also incorporate the F1-micro score, which is computed by aggregating the true positives, false positives, and false negatives across all classes before calculating the overall F1-score. This metric is particularly useful for evaluating performance in imbalanced datasets. Additionally, we include Accuracy, which is defined as the ratio of correctly classified frames to the total number of frames, to provide a more comprehensive assessment of model performance. For simplicity, we will refer to the F1-macro score as F1-score in the remainder of the text.
@@
Table 4: Quantitative comparison results on BP4D+ dataset, where the evaluation metric is F1-score(%).
Method Source AU Index Avg.
1 2 4 6 7 10 12 14 15 17 23 24
ARL [5] 2019 TAC 29.9 33.1 27.1 81.5 83.0 84.8 86.2 59.7 44.6 43.7 48.8 32.3 54.6
EAC-Net [4] 2018 TPAMI 38.0 37.5 32.6 82.0 83.4 87.1 85.1 62.1 44.5 43.6 45.0 32.8 56.1
[…]
AC2[34] 2024 IJCV 42.3 35.4 26.7 80.7 87.0 90.9 85.8 73.3 45.3 43.4 50.3 29.0 57.5
AUNCE(Ours) - 42.5 41.5 27.9 84.1 86.5 91.7 91.3 72.1 49.7 48.8 48.0 33.5 59.8
($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$
0.18) 0.46) 0.14) 0.32) 0.23) 0.17) 0.14) 0.49) 0.33) 0.25) 0.12) 0.24) 0.26)
@@
Table 7: Quantitative comparison results for common 10 AUs of cross evaluation between BP4D and GFT. BP4D → GFT denotes training on BP4D and testing on GFT. The evaluation metric is F1-score(%).
Dataset Method AU Index Avg.
1 2 4 6 10 12 14 15 23 24
BP4D → GFT AC2[34] 28.0 35.7 22.7 70.5 69.2 65.2 16.3 29.0 40.8 45.2 42.3
Ours (WCE) 30.2 37.9 24.8 66.4 72.3 62.6 20.1 28.6 44.6 38.8 42.6
Ours (AUNCE) 32.4 50.1 10.5 71.0 77.6 70.2 22.3 26.5 50.9 40.4 45.2
GFT → BP4D Ours (WCE) 49.5 45.5 26.0 20.4 52.3 36.6 17.3 18.6 50.5 54.4 37.1
AC2[34] 51.9 49.3 25.8 24.6 50.1 40.9 15.3 20.1 47.2 58.3 38.4
Ours (AUNCE) 52.3 50.8 24.6 18.7 58.8 38.8 20.6 22.5 56.6 62.9 40.7
@@ 4.2.1 Evaluation on BP4D
Methods in the first group are based on feature extraction of regional area around AUs, including EAC-Net [4], ROI [3], ARL [5] and AC2[34]. Our approach achieves a notable improvement, with an increase in F1-score ranging from 1.5% to 10.2% over these methods. This advancement is primarily due to the feature extraction methods of these frameworks neglecting critical correlation information among AUs and failing to specifically address the class imbalance issue for each AU. Furthermore, the precision of the encoders utilized by these methods is generally lower.
@@
Methods in the fourth group focus on capturing correlational information among AUs, including AAR [32], SACL [39] and MEGraph [26]. Our approach exhibits a performance advantage of approximately 0.5%-2.3% in F1-score. This underscores the efficacy of our supervised discriminative contrastive learning paradigm in effectively capturing differential information between AUs from each sample pair instead of pixel-level information of the entire face associated with AU. It is noteworthy that the encoder used in our method is only the same as the network MEGraph used for the first stage of training, with reduced parameters by nearly 7M. This highlights learning only the difference information among features is more efficient than learning the whole pixel-level information.
Methods in the last group are based on contrastive learning, including SIMCLR [13], MoCo [14], EmoCo [23], CLP [11], KSRL [21] and CLEF [19]. When compared to these methods, our approach demonstrates a performance advantage of approximately 0.2% to 16.3% in F1-score. This improvement is attributed to the efficacy of our proposed negative sample re-weighting strategy and positive sample sampling strategy, specifically tailored for the AU detection task using static images. In addition, compared to contrastive learning methods based on weighted cross-entropy loss, our approach significantly reduces the extraction of pixel-level redundant information.
@@ 4.2.2 Evaluation on DISFA
Comparable outcomes are evident in the DISFA dataset, as outlined in Table 3, where AUNCE attains the highest overall performance in AU detection. Furthermore, a comparison between Table 2 and Table 3 reveals that recent methods such as EmoCo, CLP, and MEGraph demonstrate strong performance on the BP4D dataset but achieve only mediocre results on the DISFA dataset. This disparity suggests their limited generalizability. The more pronounced class imbalance issue in DISFA, as shown in Table 1, likely contributes to this performance gap. In contrast, AUNCE consistently delivers robust performance on both BP4D and DISFA datasets and achieves more stable results across various AUs due to our proposed negative sample re-weighting strategy.
4.2.3 Evaluation on BP4D+
To assess the model’s performance on a larger and more diverse test set, we train AUNCE on the full BP4D and evaluate it on the entire BP4D+ using a cross-dataset testing protocol. The comparative results are presented in Table 4. Notably, AUNCE achieves superior performance in terms of the average F1-score, outperforming previous methods. This indicates that AUNCE maintains robust performance even when the testing data significantly increases in both scale and diversity.
@@ 4.2.4 Evaluation on GFT
The F1-score results on GFT are summarized in Table 5, where the proposed AUNCE method demonstrates a clear advantage over previous approaches. Unlike BP4D, BP4D+, and DISFA, which predominantly consist of near-frontal facial images, GFT features a substantial number of images captured with out-of-plane poses, posing greater challenges for accurate AU detection. While methods such as ARL, CLP and AC2D encounter significant performance limitations under these conditions, AUNCE achieves an impressive average F1-score of 63.0%, highlighting its robustness in handling complex pose variations.
@@ 4.2.5 Evaluation on Aff-Wild2
4.2.6 Cross dataset validation of AUNCE
To evaluate the generalization and cross-dataset reliability of our framework, we compare AUNCE with two baselines: one using the same encoder with weighted cross-entropy loss (WCE) and another, AC2[34]. Results are presented in Table 7. Since the 10 AUs in the GFT dataset overlap with the 12 AUs in BP4D, we perform cross-dataset evaluations between BP4D and GFT. For BP4D-to-GFT, we use the three BP4D models trained via 3-fold cross-validation and report the average results. For GFT-to-BP4D, the trained GFT model is tested on the three BP4D test sets, and average results are reported.
As shown in Tables 2 and 5, the performance of AUNCE is notably lower in this cross-dataset evaluation. This performance drop can be attributed to the substantial domain gap between BP4D and GFT. Additionally, our findings show that AUNCE demonstrates superior performance compared to the other two approaches that rely on WCE loss in both BP4D → GFT and GFT → BP4D tasks. This result further underscores the robustness and generalization capability of our proposed discriminative contrastive learning framework.
Table 8: Comparison of the AUNCE loss with the weighted cross-entropy loss (WCE) and the weighted focal loss (WFL) during the pretraining phase across different encoders on BP4D and DISFA datasets.
Backbone Source Loss Dataset
BP4D DISFA
ResNet50 [40] 2016CVPR WCE 56.2 45.2
WFL 57.71.5↑ 47.01.8↑
AUNCE $\textbf{59.7}^{{\color[rgb]{1,0,0}3.5\uparrow}}$ $\textbf{51.6}^{{\color[rgb]{1,0,0}6.4\uparrow}}$
ROI [3] 2017CVPR WCE 56.4 48.5
WFL 57.00.6↑ 50.62.1↑
AUNCE $\textbf{58.8}^{{\color[rgb]{1,0,0}2.4\uparrow}}$ $\textbf{52.8}^{{\color[rgb]{1,0,0}4.3\uparrow}}$